# Benchmarking Time Series Foundation Models for Short-Term Household Electricity Load Forecasting

### v1

@@ 5. Discussion
Considering the input sizes, TimesFM significantly outperforms PatchTST for longer input sizes (96 and 168). The foundation model Chronos performs comparably to PatchTST for the input sizes 96 and 168 while performing worse than the Vanilla Transformer and the PatchTST for the input size 24. This suggests that the performance of foundation models may increase with input size. A possible reason is that foundation models need more context than custom-trained Transformers to identify the patterns of the time series.
A potential for the good performance of TimesFM could be that its training data contains one household energy dataset (Das et al., 2024), while Chronos’ training set does not comprise any energy time series (Ansari et al., 2024). It is conceivable that therefore TimesFM recognized certain patterns in the household energy load better than Chronos. This observation is a hint that researchers and practitioners should pay careful attention to the respective training data when choosing an FM model.
@@
LagLlama performs poorly for all input sizes compared to the other foundation models and most TFS Transformers. This may be explained by LagLlama’s architecture, which uses lags to predict future values (Rasul et al., 2024). These lags include quarterly, monthly, weekly, daily, and hourly levels (Rasul et al., 2024), while our defined maximum input size of 168 hours allows only incorporating daily to weekly lags.
@@
Considering the performance of the TFS Transformer PatchTST, our study supports the findings of Hertel et al. (2023) and Cen and Lim (2024), as the PatchTST model provides the second best overall performance and even slightly outperforms TimesFM on an input size of 24 hours. Additionally, the results of Hertel et al. (2023) indicate that TFS Transformers perform better when trained globally with more data. Our study extends their findings by showing that the zero-shot performance of the also globally trained foundation models TimesFM and Chronos is better or comparable to TFS Transformers.
@@
A final observation of our analysis is that the SeasonalAverage baseline outperforms some of the TFS Transformers and LagLlama for horizons of 96 hours and 168 hours. This is probably due to the strong daily patterns present in household energy load.
6. Limitations & Outlook
@@
Naturally, our analysis is not without limitations. Hyperparameter tuning the TFS Transformers may increase the performance of these models supporting the findings of Sievers and Blank (2023) and Upadhyay et al. (2023) while simultaneously increasing the training effort compared to foundation models even more. On the other hand, longer input sizes could lead to better performance of the foundation models, especially for LagLlama, which needs higher context lengths to generate the lagged features (Rasul et al., 2024).
@@
There are several directions for future research on household STLF. First, incorporating longer input sizes would allow for drawing a better picture of TFS Transformer-based and foundation model behavior, e.g., the behavior of LagLlama on longer inputs. While foundation models provide zero-shot forecasts that can outperform SOTA TFS Transformers, fine-tuning foundation models has been shown to increase performance in other disciplines, such as foundation LLMs. Moreover the foundation model TimesFM, which has been trained on some household energy data, performed better than other FMs. Hence, pre-training and fine-tuning foundation models on energy time series forecasting could further increase their capabilities. Second, we propose that the models can learn cross-domain global patterns, such as the COVID-19 pandemic or geopolitical crises, which may lead to information leakage when evaluating these models on holdout datasets originating from the same time period as the training datasets. This possible problem should be explored in future research. Third, our univariate analysis could be extended to multivariate foundation models, including covariates such as weather data (Kang and Reiner, 2022).
7. Conclusion
@@
Motivated by recent algorithmic developments in time series forecasting, this study, as the first in household electricity STLF, investigated whether time series foundation models are competitive to SOTA TFS Transformers on household STLF tasks. Following the guideline of Haben et al. (2021) we considered time series cross-validation on multiple datasets using statistical baselines and sophisticated TFS Transformers. In our benchmark, unlike the TFS Transformers, the foundation models were used out-of-the-box for prediction without task-specific adaptation or fine-tuning. We show that the foundation models are already capable of delivering competitive forecast performance compared to trained-from-scratch Transformers. The foundation models show their strength, especially when there is more context (i.e., longer input sizes). Furthermore, our findings indicate that, in the case of LagLLama, its special architecture may harm performance when dealing with a limited input context. In conclusion, the ability of foundation models to achieve high accuracy with limited data and without training opens up new possibilities for developing more efficient and accessible energy forecasting solutions.

### v2

@@ 5 Discussion
Going more into detail and considering the input sizes, TSFM like Sundial, TimesFM or Chronos-Bolt significantly outperform PatchTST for longer input sizes (96 and 168) in regards of MAE. This suggests that the performance of foundation models may increase with input size. A possible reason is that TSFM need more context than custom-trained Transformers to identify the patterns of the time series.
@@
LagLlama performs poorly for all input sizes compared to the other foundation models and most TFS Transformers. This might be explained by LagLlama’s architecture, which uses lags to predict future values Rasul et al. (2024). These lags include quarterly, monthly, weekly, daily, and hourly levels Rasul et al. (2024), while our defined maximum input size of 168 hours allows only incorporating daily to weekly lags. Similarly our restricted experimental setup with limited information about the time series could also explain the performance of the Moirai model, which was originally evaluated with a significantly longer context size of 1000 Woo et al. (2024).
@@
Considering the performance of the TFS Transformer PatchTST, our study supports the findings of Hertel et al. (2023) and Cen and Lim (2024), as the PatchTST model provides good overall performance and even slightly outperforms other TSFM on an input size of 24 hours. Additionally, the results of Hertel et al. (2023) indicate that TFS Transformers perform better when trained globally with more data. Our study extends their findings by showing that the zero-shot performance of the also globally trained TSFM is better or comparable to TFS Transformers. Interestingly, we could not reproduce the statement that input sizes exceeding the horizon decreases forecasting performance of transformer models Wen et al. (2022), but observed the opposite: PatchTST as well as all TSFM benefit from longer context. Just the the VanillaTransformer did not benefit from longer input sizes matching the results of Wen et al. (2022).
Moreover, when assessed using the household STLF-specific metric (APNEh), TSFM models like Time-MoE and Sundial are able to deliver good results, but the difference to the baseline is significantly smaller. Most TSFM seem to predict more conservative without extreme load peaks. As these peaks are also relevant in STLF, e.g. regarding grid capacity and stability, TSFM seem not to be suitable for predicting these load peaks.
@@
A final observation of our analysis is that the SeasonalAverage baseline outperforms some of the TFS Transformers and TSFM for horizons of 96 hours and 168 hours. This is probably due to the strong daily patterns present in household energy load, which makes the SeasonalAverage a suitable baseline.
Compared to standard TSFM evaluations (see Section 2), our approach combines multi-dataset evaluation with time-series cross-validation for a more comprehensive assessment. This design mitigates the limitations of smaller datasets by capturing both cross-series variation and performance changes over time, thereby strengthening the validity of our findings and supporting stronger claims about the robustness and generalizability of TSFMs relative to the TFS Transformer.
6 Limitations & Outlook
@@
Naturally, our analysis is not without limitations. Hyperparameter tuning the TFS Transformers may increase the performance of these models supporting the findings of Sievers and Blank (2023) and Upadhyay et al. (2023) while simultaneously increasing the training effort compared to TSFM even more. On the other hand, longer input sizes could lead to better performance of the foundation models, especially for LagLlama and Moirai, which might need higher context lengths due to their models architecture Rasul et al. (2024); Woo et al. (2024).
@@
There are several directions for future research on household STLF. First, incorporating longer input sizes would allow for drawing a better picture of TFS Transformer-based and foundation model behavior, e.g., the behavior of LagLlama on longer inputs. While TSFM provide zero-shot forecasts that can outperform SOTA TFS Transformers, fine-tuning foundation models has been shown to increase performance in other disciplines, such as foundation LLMs. Moreover TSFM, which have been trained on some household energy data, performed better than other TSFMs. Hence, pre-training and fine-tuning foundation models on energy time series forecasting could further increase their capabilities. Second, we propose that the models can learn cross-domain global patterns, such as the COVID-19 pandemic or geopolitical crises, which may lead to information leakage when evaluating these models on holdout datasets originating from the same time period as the training datasets. This possible problem should be explored in future research. Third, our univariate analysis could be extended to multivariate TSFMs, including covariates such as weather data Kang and Reiner (2022).
7 Conclusion
@@
Motivated by recent algorithmic developments in time series forecasting, this study investigated whether time series foundation models are competitive to SOTA TFS Transformers on household STLF tasks. Following the guideline of Haben et al. (2021) we considered time series cross-validation on multiple datasets using statistical baselines and sophisticated TFS Transformers. In our benchmark, unlike the TFS Transformers, the TSFMs were used out-of-the-box for prediction without task-specific adaptation or fine-tuning. We show that TSFM are already capable of delivering competitive and in most cases better forecast performance compared to trained-from-scratch Transformers. The foundation models show their strength, especially when there is more context (i.e., longer input sizes). Furthermore, our findings indicate that, in the case of LagLLama or Moirai, its special architecture may harm performance when dealing with a limited input context. In conclusion, the ability of foundation models to achieve high accuracy with limited data and without training opens up new possibilities for developing more efficient and accessible energy forecasting solutions.